# What Matters for Bioacoustic Encoding

### v1

@@ 1 Introduction
However, large amounts of unannotated or weakly labeled bioacoustic data are recorded regularly, especially through Passive Acoustic Monitoring (PAM) (Gibb et al., 2019), and citizen science platforms such as Xeno-Canto (Vellinga & Planqué, 2015) or iNaturalist (Chasmai et al., 2024). These data can be leveraged to train a bioacoustic encoder, which can then be deployed in downstream tasks, as bioacoustic features (e.g. for linear probing or clustering) or finetuning the whole model, among other options. Two such classic and state-of-the-art bioacoustic encoders are BirdNet (Kahl et al., 2021) and Perch (Ghani et al., 2023) which have been applied to downstream application tasks such as multi-taxa species retrieval and detection (Pérez-Granados, 2023; Dumoulin et al., 2025; Ghani et al., 2023)
Other bioacoustic encoders have been proposed, to be reviewed in the next section. They have in common a supervised learning approach, usually limited to a single taxonomic group, with notable exceptions including SurfPerch (Williams et al., 2024) and recently the models of iNatSounds (Chasmai et al., 2025). Moreover, they evaluate the quality of the learned representations on a limited set of downstream tasks and datasets. Typically they solely evaluate on species classification, with their training and test data containing the same species, often with an out-of-distribution effect, as training datasets typically consists of focal recordings, while evaluation datasets are soundscape recordings (Rauch et al., 2025b). In contrast, real-world bioacoustic applications require encoders that generalize effectively across diverse species and tasks, often beyond those explicitly seen during training. For example, researchers may need to recognize previously unobserved species, identify individual animals from limited vocal data, or characterize animal vocal repertoires without extensive annotations. Evaluating models on such diverse and realistic scenarios is critical, yet building and measuring performance of encoders that generalize across these conditions remains underexplored in current works.
@@
We find that under comparable training conditions, self-supervised models achieve strong out-of-distribution generalization yet under-perform supervised models on in-distribution tasks, and that incorporating general audio into bioacoustic training significantly improves model transferability. Leveraging these insights, we propose a set of training recipes and models that achieve state-of-the-art results overall on our extensive evaluation benchmark, offering a versatile encoder for bioacoustic research. In contrast to previous works (Ghani et al., 2025; Kather et al., 2025) we find that combining self-supervised and supervised learning on diverse data, i.e. including general audio at pre-training and post-training, brings benefits when transferring these representations to detection and classification benchmarks.
@@ 2 Related Work
An extensive number of works propose audio and speech encoders, most of them transformer-based, such as Wav2vec (Baevski et al., 2020), HuBERT (Hsu et al., 2021), AudioMAE (Huang et al., 2022), BEATs (Chen et al., 2023) or EAT (Chen et al., 2024). Several bioacoustic-specific encoders have also been developed: BirdNet (Kahl et al., 2021) and Perch (Ghani et al., 2023) build upon an EfficientNet (EffNet) architecture (Tan & Le, 2021), a CNN-based vision neural network pre-trained on ImageNet (Russakovsky et al., 2015) taking audio spectrograms as input. Transformer-based bioacoustic models include AVES (Hagiwara, 2023) based on HuBERT, Animal2Vec (Schäfer-Zimmermann et al., 2024) based on data2vec (Baevski et al., 2023), BirdMAE (Rauch et al., 2025a) based on AudioMAE with modified decoder architecture, and TweetyBert (Vengrovski et al., 2025) inspired by BERT (Devlin et al., 2019). However, these approaches lack systematic comparison across different architectures with standardized pipelines. In contrast, we compare with a wide range of encoder baselines and aim to have a fair comparison across different architectures, with minimal changes and near-identical pipelines.
@@
Bioacoustic encoders have been evaluated primarily in the context of species classification and detection (Rauch et al., 2025b; Hamer et al., 2023; Ghani et al., 2023; Chasmai et al., 2025; Kather et al., 2025) while other important tasks such as vocal repertoire discovery(Anikin et al., 2018) or individual identification(Stowell et al., 2019) have been scarcely addressed. These tasks, which are critical to the study of animal communication yet lack large-scale annotated data, are a natural test-bed for generalization of learned representations. Research on these two topics has so far used a limited number of private datasets or has not compared with state-of-the-art bioacoustic encoders (Best et al., 2023; Nolasco et al., 2025; Stowell et al., 2019; Wierucka et al., 2025). To give a broader overview of the capabilities of a bioacoustic encoder, we address these limitations by adding 8 public datasets, not considered in any previous benchmark. Further, related work from Kather et al. (2025) and from Best et al. (2023) has gained insight into bioacoustic encoders by analyzing their embeddings with clustering metrics, including a qualitative finding that self-supervised encoders better generalized from birds to frogs. We introduce a similar evaluation methodology, enhancing BEANS (Hagiwara et al., 2023) and BirdSet (Kather et al., 2025) with clustering and retrieval metrics, scaling a related analysis from two datasets to twenty-six. For a comparison with current bioacoustics benchmarks, we present Table 1, which includes training and evaluation configurations.
@@ 3.4 Evaluation Setup: Data, Tasks, and Metrics
For all tasks other than Vocal Repertoire Discovery we report classic performance metrics on linear probing (top-1 accuracy for single-label tasks, macro-averaged mean average precision for multilabel tasks). To evaluate retrieval, we consider each item in the test set as a query. We rank the remaining elements of the test set according to their cosine similarity with the query under the model’s embedding function, and evaluate the ordering. For single-label tasks (BEANS Classification, Individual ID, and Vocal Repertoire), we consider items as relevant if they share the same label as the query. For multi-label tasks (BEANS Detection, BirdSet) we consider a query to be correct it shares at least one label with the ground-truth item. We exclude queries with no labels, while ground-truth items with no labels are included in the evaluation as negatives. We measure this ranking using ROC AUC as the primary metric (marked as R-AUC with R for retrieval). For Individual Identification evaluation datasets Pipit and Chiffchaff (Stowell et al., 2019), the training set follows an intentionally different background noise distribution than the test set following original design in Stowell et al. (2019). To evaluate generalization in this scenario, we report ROC AUC by using each example in the train set as a query, rather than each example in the test set as a query. For single-label datasets in all benchmarks, we also consider a clustering task. For each evaluation set, we perform K-means clustering given a known number of clusters (labels). We evaluate similarity of these clusters compared to the known class groups by measuring normalized mutual information (marked as NMI) between the two. These two additional evaluations offer insight into the pre-trained embeddings of a model, but are also relevant to downstream tasks in bioacoustics, namely audio-to-audio retrieval (Hamer et al., 2023) (retrieval) and repertoire discovery (Best et al., 2023) (clustering) or call-type classification. For vocal repertoire discovery, we use clustering and retrieval as the primary evaluations. All metrics are formalized in the Appendix in A.1.
@@ 4 Results
4.1 Post-training EfficientNet
@@
4.2 Self-supervised Pre-training
@@
4.3 Post-training Self-Supervised Backbones

### v2

@@ 1 Introduction
However, large amounts of unannotated or weakly labeled bioacoustic data are recorded regularly, especially through Passive Acoustic Monitoring (PAM) (Gibb et al., 2019), and citizen science platforms such as Xeno-Canto (Vellinga & Planqué, 2015) or iNaturalist (Chasmai et al., 2024). These data can be leveraged to train a bioacoustic encoder, which can then be deployed in downstream tasks, as bioacoustic features (e.g. for linear probing or clustering) or finetuning the whole model, among other options. Two such classic and state-of-the-art bioacoustic encoders are BirdNet (Kahl et al., 2021) and Perch (Ghani et al., 2023) which have been applied to downstream application tasks such as multi-taxa species retrieval and detection (Pérez-Granados, 2023; Dumoulin et al., 2025; Ghani et al., 2023).
Other bioacoustic encoders have been proposed, to be reviewed in the next section. They have in common a supervised learning approach, usually limited to a single taxonomic group, with notable exceptions including SurfPerch (Williams et al., 2024) and recently the models of iNatSounds (Chasmai et al., 2025). Moreover, they evaluate the quality of the learned representations on a limited set of downstream tasks and datasets. Typically they solely evaluate on species classification, with their training and test data containing the same species, often with an out-of-distribution effect, as training datasets typically consists of focal recordings, while evaluation datasets are soundscape recordings (Rauch et al., 2025b). In contrast, real-world bioacoustic applications require encoders that generalize effectively across diverse species and tasks, often beyond those explicitly seen during training. For example, researchers may need to recognize previously unobserved species, identify individual animals from limited vocal data, or characterize animal vocal repertoires without extensive annotations. Evaluating models on such diverse and realistic scenarios is critical, yet building and measuring the performance of encoders that generalize across these conditions remains underexplored in current works.
@@
We find that under comparable training conditions, self-supervised models achieve strong out-of-distribution generalization yet under-perform supervised models on in-distribution tasks, and that incorporating general audio into bioacoustic training significantly improves model transferability. Sequential self-supervised and supervised learning yields strong performance both in and out-of-distribution. Leveraging these insights, we propose a set of training recipes and models that achieve state-of-the-art results overall on our extensive evaluation benchmark, offering a versatile encoder for bioacoustic research.
@@ 2 Related Work
An extensive number of works propose audio and speech encoders, most of them transformer-based, such as Wav2vec (Baevski et al., 2020), HuBERT (Hsu et al., 2021), AudioMAE (Huang et al., 2022), BEATs (Chen et al., 2023) or EAT (Chen et al., 2024). Several bioacoustic-specific encoders have also been developed: BirdNet (Kahl et al., 2021) and Perch (Ghani et al., 2023) build upon an EfficientNet (EffNet) architecture (Tan & Le, 2021), a CNN-based vision neural network pre-trained on ImageNet (Russakovsky et al., 2015) taking audio spectrograms as input. Transformer-based bioacoustic models include AVES (Hagiwara, 2023) based on HuBERT, Animal2Vec (Schäfer-Zimmermann et al., 2024) based on data2vec (Baevski et al., 2023), BirdMAE (Rauch et al., 2025a) based on AudioMAE with modified decoder architecture, and TweetyBert (Vengrovski et al., 2025) inspired by BERT (Devlin et al., 2019). However, within bioacoustics these approaches lack systematic comparison across different architectures with standardized pipelines. In contrast, we compare with a wide range of encoder baselines and aim to have a fair comparison across different architectures, with minimal changes and near-identical pipelines.
@@
Bioacoustic encoders have been evaluated primarily in the context of species classification and detection (Rauch et al., 2025b; Hamer et al., 2023; Ghani et al., 2023; Chasmai et al., 2025; Kather et al., 2025) while other important tasks such as vocal repertoire discovery(Anikin et al., 2018) or individual identification(Stowell et al., 2019) have been scarcely addressed. These tasks, which are critical to the study of animal communication yet lack large-scale annotated data, are a natural test-bed for generalization of learned representations. Research on these two topics has so far used a limited number of private datasets or has not compared with state-of-the-art bioacoustic encoders (Best et al., 2023; Nolasco et al., 2025; Stowell et al., 2019; Wierucka et al., 2025). To give a broader overview of the capabilities of a bioacoustic encoder, we address these limitations by adding 8 public datasets, not considered in any previous benchmark. Further, related work from Kather et al. (2025) and from Best et al. (2023) has gained insight into bioacoustic encoders by analyzing their embeddings with clustering metrics, including a qualitative finding that self-supervised encoders better generalized from birds to frogs. We introduce a similar evaluation methodology, enhancing BEANS (Hagiwara et al., 2023) and BirdSet (Rauch et al., 2025b) with clustering and retrieval metrics, scaling a related analysis from two datasets to twenty-six. For a comparison with current bioacoustics benchmarks, we present Table 1, which includes training and evaluation configurations.
@@ 3.4 Evaluation Setup: Data, Tasks, and Metrics
For all tasks other than Vocal Repertoire Discovery we report classic performance metrics on linear probing (top-1 accuracy for single-label tasks, macro-averaged mean average precision for multilabel tasks). To evaluate retrieval, we consider each item in the test set as a query. We rank the remaining elements of the test set according to their cosine similarity with the query under the model’s embedding function, and evaluate the ordering. For single-label tasks (BEANS Classification, Individual ID, and Vocal Repertoire), we consider items as relevant to the query if they share the same label as the query. For multi-label tasks (BEANS Detection, BirdSet) we consider items to be relevant if they share at least one label with the ground-truth item. We exclude queries with no labels, while ground-truth items with no labels are included in the evaluation as negatives. We measure this ranking using ROC AUC as the primary metric (marked as R-AUC with R for retrieval). For Individual Identification evaluation datasets Pipit and Chiffchaff (Stowell et al., 2019), the training set follows an intentionally different background noise distribution than the test set following original design in Stowell et al. (2019). To evaluate generalization in this scenario, we report ROC AUC by using each example in the train set as a query, rather than each example in the test set as a query. For single-label datasets in all benchmarks, we also consider a clustering task. For each evaluation set, we perform K-means clustering given a known number of clusters (labels). We evaluate similarity of these clusters compared to the known class groups by measuring normalized mutual information (marked as NMI) between the two. These two additional evaluations offer insight into the pre-trained embeddings of a model, but are also relevant to downstream tasks in bioacoustics, namely audio-to-audio retrieval (Hamer et al., 2023) (retrieval) and repertoire discovery (Best et al., 2023) (clustering) or call-type classification. For vocal repertoire discovery, we use clustering and retrieval as the primary evaluations. All metrics are formalized in the Appendix in A.1.
@@ 4 Results
4.1 What Data Matters in Post-training EfficientNet
@@
4.2 Self-supervised Pre-training Helps Out-of-distribution
@@
4.3 Post-training Recipes for Self-Supervised Backbones